what is the ninth term when the first term is 4 and the common ratio is -6?

Answer by stanbon(75887)

what is the ninth term when the first term is 4 and the common ratio is -6?
----
Form: a(n) = a*r^(n-1)
---
9th term: a(9) = 4*(-6)^(8) = 6,718,464
---
Cheers,
Stan H.
====